### v1

@@ Abstract
Formal analysis of vision-based deep neural networks (DNNs) is highly desirable but it is very challenging due to the difficulty of expressing formal specifications for vision tasks and the lack of efficient verification procedures. In this paper, we propose to leverage emerging multimodal, vision-language, foundation models (VLMs) as a lens through which we can reason about vision models. VLMs have been trained on a large body of images accompanied by their textual description, and are thus implicitly aware of high-level, human-understandable concepts describing the images. We describe a logical specification language $\texttt{Con}_{\texttt{spec}}$ designed to facilitate writing specifications in terms of these concepts. To define and formally check $\texttt{Con}_{\texttt{spec}}$ specifications, we leverage a VLM, which provides a means to encode and efficiently check natural-language properties of vision models. We demonstrate our techniques on a ResNet-based classifier trained on the RIVAL-10 dataset leveraging CLIP as the multimodal model.
Keywords:
Formal analysis Concepts Vision-language models.
@@ 1 Introduction
Deep neural networks (DNNs) are increasingly used in safety-critical systems as perception components processing high-dimensional image data [13, 20, 22, 3]. Formal analysis of these networks is highly desirable but it is very challenging due to the difficulty of expressing formal specifications about vision-based DNNs. This in turn is due to the low-level nature of pixel-based input representations that these models operate on and to the fact that DNNs are also notoriously opaque—their internal computational structures remain largely uninterpreted. There are also serious scalability issues that impede formal, exhaustive verification: these networks are very large, with thousands or million of parameters, making the verification problem very complex.
@@
To demonstrate our techniques for analyzing vision-based DNNs, we consider a performant and complex ResNet-based classifier trained on the RIVAL10 dataset [24]. RIVAL10 is a subset of ImageNet [9], restricted to 10 classes. Importantly, RIVAL10 images come with manually annotated attributes which we leverage to validate our techniques. We use the popular vision-language CLIP model as the VLM in our analysis.
@@ 4 Vision-Language Models as Analysis Tools
The parametric nature of the $\texttt{Con}_{\texttt{spec}}$ semantics with respect to the concept representations map $rep$ necessitates giving a concrete implementation of $rep$ before we can verify if a model satisfies a $\texttt{Con}_{\texttt{spec}}$ specification. The map $rep$ needs to be semantically meaningful—ideally, the strength of a concept in an input as per $rep$ should be in congruence with the ground truth. For the purpose of verification, it is also essential to ensure that $rep$ can be encoded as efficiently checkable constraints. In order to implement $rep$, we could leverage an empirical observation from recent work that a neural classifier $f:X\rightarrow Y$ with encoder $f_{enc}:X\rightarrow Z$ and head $f_{head}:Z\rightarrow Y$ learns to represent concepts as directions in its representation space $Z$ [38, 23, 36]. Consequently, the strength with which an input $x$ contains a concept $con$ is proportional to the alignment between the embedding $f_{enc}(x)$ of $x$ and the direction corresponding to $con$. Defn. 2 formalizes the notion of a concept representation map implemented in this manner. Alignment with the concept is measured via the cosine similarity between the input embedding and a vector in the direction of the concept.
Definition 2 (Linear $rep$ via vision model)
Given a vision model $f:X\rightarrow Y$ that can be decomposed into $f_{enc}:X\rightarrow Z$ and $f_{head}:Z\rightarrow Y$ where $Z:=\mathbb{R}^{d^{\prime}}$, the linear concept representation map, $rep$, via model $f$ is defined as,
$rep(con):=\lambda x.\cos(f_{enc}(x),\overline{con})$
where $\overline{con}$ refers to a vector in $Z$ whose direction corresponds to the concept $con$. We assume that every concept $con\in Concepts$ is represented by a direction in $Z$ and these directions are pre-computed.
In practice, inferring the directions in the representation space of a vision model corresponding to concepts can be challenging. First, the hypothesis that concepts are represented as directions in embedding space may itself be invalid as has been suggested in some other works [7, 1]. Even if it holds, finding the direction corresponding to a concept requires data annotated with concepts and such data need not be available.
@@
In order to tackle these challenges we propose the use of vision-language models (VLMs) such as CLIP [28] as a means for implementing $rep$. Using VLMs to implement $rep$ is advantageous for several reasons. First, the manner in which VLMs such as CLIP are trained ensures that concepts are represented as directions in the common VLM representation space. In particular, their training objective requires the pair of embeddings generated for a image-caption pair, by the image and text encoders of a VLM, to be aligned in the same direction as measured via cosine similarity. Second, since the two encoders map to the same representation space, it allows the use of language for probing and reasoning about visual data. For instance, one can simply compute the embedding of a caption which references the relevant concept (say, “Image of a wheel” in case of our running example) to obtain the concept direction. This obviates the need to have image data annotated with concepts. Third, VLMs are typically treated as foundation models [5] that are well-trained on vast and diverse datasets to serve as building blocks for downstream applications. As a result, these models are exposed to a variety of concepts from multiple datasets and can therefore serve as a useful repository of concept representations. The richness of the VLM embedding space also opens up the opportunity to define aggregate concepts such as “wheels and metallic” where the logical connectives are incorporated into the concept itself instead of expressing them externally via the connectives of $\texttt{Con}_{\texttt{spec}}$, thereby simplifying the structure of a $\texttt{Con}_{\texttt{spec}}$ specification at the cost of more complex concepts. We plan to explore this direction more in the future.
@@
Defn. 3 describes the implementation that we use for the concept representation map $rep$ that uses a VLM $g$. While the definition still relies on the hypothesis that concepts are represented as directions, the direction corresponding to a concept $con$ is extracted using the text encoder $g_{enc}^{txt}$ of $g$. In particular, following existing work [25], a vector $\overline{con}$ in the direction of concept is computed as the mean of text embeddings for a set of captions (denoted as $cap(con)$) that all refer to $con$ in different ways. Taking the mean is essential since many different valid captions can be constructed for the same concept and each caption leads to a slightly different embedding. Notice that there is no need for images with concept annotations to infer the concept directions.
@@
Given a VLM $g$ with image encoder $f_{enc}^{img}:X\rightarrow Z$ and text encoder $f_{enc}^{txt}:T\rightarrow Z$ where $Z:=\mathbb{R}^{d^{\prime}}$, the linear concept representation map, $rep$, via model $g$ is defined as,
@@
The advantages of VLM-based concept representations as defined in Defn. 3 make them a good choice for verifying models with respect to $\texttt{Con}_{\texttt{spec}}$ specifications. However, they also introduce a major challenge—in the course of verifying a vision model $f$, we also need to reason about the complex image encoder $g_{enc}^{img}$ of a VLM $g$ since $g_{enc}^{img}$ is used to implement the concept representation map. Fortunately, it has been recently observed [25] that the representation space of vision-based DNNs and VLMs such as CLIP can be linked to each other via a affine map, i.e., an embedding of an input image in the representation space of a vision-based DNN can be mapped to the corresponding image embedding in the representation space of a VLM via a affine map. As a consequence, assuming that the map causes no loss of information, it becomes possible to use a VLM-based concept representation map without referring to $g_{enc}^{img}$.
Formally, let $D_{\text{train}}\subset X$ denote the training dataset. Given the encoder $f_{enc}:X\rightarrow Z_{f}$ of a vision model $f$ and the image encoder $g_{enc}^{img}:X\rightarrow Z_{g}$ of VLM $g$, representation space alignment of $f_{enc}$ to $g_{enc}^{img}$ is the task of learning a mapping $r_{map}:Z_{f}\rightarrow Z_{g}$. Moayeri et al. [25] show that $r_{map}$ can be restricted to the class of affine transformations, i.e., $r_{map}(z):=Mz+d$, where $r_{map}$ is learnt by solving the following optimization problem,
@@
If the representation space of the encoder $f_{enc}:X\rightarrow Z_{f}$ of a vision model $f$ is mapped to the representation space of the image encoder $g_{enc}^{img}:X\rightarrow Z_{g}$ of a VLM $g$ through the map $r_{map}$, then Defn. 4 shows that the VLM-based concept representation map (defined in Defn. 3) can be expressed using only $f_{enc}$ and $r_{map}$. This, in turn, ensures that we do not need to encode the complex VLM image encoder when verifying the vision model $f$ with respect to a $\texttt{Con}_{\texttt{spec}}$ specification.
Definition 4 (Linear $rep$ via vision model and $r_{map}$ )
Given a vision model $f:X\rightarrow Y$ with encoder $f_{enc}:X\rightarrow Z_{f}$, and a VLM $g$ with encoders $g_{enc}^{img}:X\rightarrow Z_{g}$ and $g_{enc}^{txt}:T\rightarrow Z_{g}$, if the representation space of $f_{enc}$ is mapped to the representation space of $g_{enc}^{img}$ via linear map $r_{map}$ then the linear concept representation map, $rep$, can be defined as,
@@
where $\overline{con}$ is a vector in $Z_{g}$ whose direction corresponds to concept $con$.
@@
The ability to decompose a vision model $f$ into an encoder $f_{enc}$ and a $f_{head}$ can be exploited to simplify verification of vision models with respect to a $\texttt{Con}_{\texttt{spec}}$ specification. In particular, Thm. 4.1 shows that $f$ satisfies a $\texttt{Con}_{\texttt{spec}}$ specification $e$ with respect to a concept representation map $rep$ (where $rep$ is implemented as in Defn. 2 or 4) and an input scope $B$ if and only if it is also the case that $f_{head}$ also satisfies $e$ with respect to a version of $rep$, denoted as $\widehat{rep}$ modified to operate on embeddings instead of inputs (by replacing all appearances of $f_{enc}$ in Defn. 2 or 4 with the identity function) and the image of set $B$ under $f_{enc}$. As a result of this theorem, one only needs to prove that $f_{head}$ satisfies the specification and can avoid the need to translate $f_{enc}$ into checkable constraints. Given the fact that, often, the encoder of a model is much more complicated than the head (for instance, the head might only comprise the last one or two layers of the model), this can lead to significant simplification of the verification problem to be solved.
@@
Given a vision model $f:X\rightarrow Y$ that can be decomposed into $f_{enc}:X\rightarrow Z$ and $f_{head}:Z\rightarrow Y$ where $Z:=\mathbb{R}^{d^{\prime}}$, a $\texttt{Con}_{\texttt{spec}}$ specification $e$, a linear concept representation map $rep$ (as defined in Defn. 2 or 4), and an input scope $B\subseteq X$,
@@
where $\widehat{rep}$ is obtained from $rep$ by replacing $f_{enc}$ with the identity function and $f_{enc}(B):=\{f_{enc}(x)\mid x\in B\}$.
@@
See Appendix 0.B.
@@ 5 Case Study
5.1 Dataset, Concepts, and Stregth Predicates
@@
[24] is a dataset consisting of 26k images drawn from ImageNet [9], organized into 10 classes matching those of CIFAR10. The dataset contains manually annotated instance-wise labels for 18 attributes (which we use to define set $Concepts$) as well as the respective segmentation masks for these attributes on the images.
@@
The information in Figure 4 can be further used to elicit strength predicates relevant for each class $c$—for each pair of relevant ($con_{r}$) and irrelevant ($con_{ir}$) concepts, we can define a predicate $con_{r}\texttt{>}con_{ir}$ that we expect to hold on inputs of class $c$. For instance, for class truck, we formulate a total of 72 predicates, corresponding to all the possible combinations of relevant and irrelevant concepts; e.g., we elicit $wheels\texttt{>}wings$ and $metallic\texttt{>}tail$, as $wheels$ and $metallic$ are relevant concepts, whereas $wings$ and $tail$ are not. Overall, we formulate a total of 522 such predicates for the model.
@@ 5.3 Extraction of Concept Representations
In order to analyze (test or verify) the models with respect to the specifications that we formulate for RIVAL10, we need to build the concept representation map $rep$ using the VLM (as defined in Defn. 4). In particular, we need to extract the directions corresponding to the relevant concepts in CLIP’s representation space and learn the affine map from the representation space of ResNet18 to CLIP. The first step is to define the set of relevant concepts. For RIVAL10, these are the 18 attributes defined by the developers of the datatset. In general, the relevant concepts are elicited in collaboration with domain experts.
Next, to extract the directions corresponding to these 18 concepts, we use CLIP’s text encoder (referred to as $g_{enc}^{txt}$) in a manner similar to the approach described by Moayeri et al. [25]. In particular, as described in Defn. 3, for each concept $con$, we create a set of captions that refer to the concept (denoted as $cap(con)$). As an example, for the concept metallic, we use captions such as “a photo containing a metallic object”, “a photo of a metallic object”, etc. The complete set of captions used is given in Appendix 0.C. We then apply $g_{enc}^{txt}$ to each caption in the set $cap(con)$ and compute the mean of the resulting embeddings. The direction of the resulting mean vector corresponds to $\overline{con}$, i.e., the direction representing the concept $con$ in CLIP’s representation space.
The final step is to learn the affine map $r_{map}$ from the representation space of the ResNet18 vision model to the representation space of the CLIP model used in our case-study by solving the optimization problem in Equation 1. The problem is solved via gradient descent, following the approach of Moayeri et al. [25], on RIVAL10 training data for $50$ epochs using a learning rate of $0.01$, momentum of $0.9$, and weight decay of $0.0005$. The learnt map, of type $\mathbb{R}^{512}\rightarrow\mathbb{R}^{512}$, has a low Mean Squared Error (MSE) of 0.963 and a high Coefficient of Determination (R2) of 0.786 on the test data, suggesting that it is able to align the two spaces.

### v3

@@ Abstract
The analysis of vision-based deep neural networks (DNNs) is highly desirable but it is very challenging due to the difficulty of expressing formal specifications for vision tasks and the lack of efficient verification procedures. In this paper, we propose to leverage emerging multimodal, vision-language, foundation models (VLMs) as a lens through which we can reason about vision models. VLMs have been trained on a large body of images accompanied by their textual description, and are thus implicitly aware of high-level, human-understandable concepts describing the images. We describe a logical specification language $\texttt{Con}_{\texttt{spec}}$ designed to facilitate writing specifications in terms of these concepts. To define and formally check $\texttt{Con}_{\texttt{spec}}$ specifications, we build a map between the internal representations of a given vision model and a VLM, leading to an efficient verification procedure of natural-language properties for vision models. We demonstrate our techniques on a ResNet-based classifier trained on the RIVAL-10 dataset using CLIP as the multimodal model.
@@ 1 Introduction
Deep neural networks (DNNs) are increasingly used in safety-critical systems as perception components processing high-dimensional image data [13, 20, 22, 3]. The analysis of these networks is highly desirable but it is very challenging due to the difficulty of expressing formal specifications about vision-based DNNs. This in turn is due to the low-level nature of pixel-based input representations that these models operate on and to the fact that DNNs are also notoriously opaque—their internal computational structures remain largely uninterpreted. There are also serious scalability issues that impede formal, exhaustive verification: these networks are very large, with thousands or million of parameters, making the verification problem very complex.
@@
To demonstrate our techniques for analyzing vision-based DNNs, we consider a performant and complex ResNet-based classifier trained on the RIVAL10 dataset [24]. RIVAL10 is a subset of ImageNet [9], restricted to 10 classes. Importantly, RIVAL10 images come with annotated attributes which we leverage to define candidate properties to verify. In practice, we expect domain experts to provide the properties, similar to all work on verification. We use the popular vision-language CLIP model as the VLM in our analysis.
@@ 4 Vision-Language Models as Analysis Tools
The parametric nature of the $\texttt{Con}_{\texttt{spec}}$ semantics with respect to the concept representations map $rep$ necessitates giving a concrete implementation of $rep$ before we can verify if a model satisfies a $\texttt{Con}_{\texttt{spec}}$ specification. The map $rep$ needs to be semantically meaningful—ideally, the strength of a concept in an input as per $rep$ should be in congruence with the human understanding. For the purpose of verification, it is also essential to ensure that $rep$ can be encoded as efficiently checkable constraints.
@@
We propose the use of vision-language models (VLMs) such as CLIP [28] as a means for implementing $rep$. VLMs are typically treated as foundation models [5] that are well-trained on vast and diverse datasets to serve as building blocks for downstream applications. As a result, these models are exposed to a variety of concepts from multiple datasets and can therefore serve as a useful repository of concept representations. The richness of the VLM embedding space also opens up the opportunity to define aggregate concepts such as “wheels and metallic” where the logical connectives are incorporated into the concept itself instead of expressing them externally via the connectives of $\texttt{Con}_{\texttt{spec}}$, thereby simplifying the structure of a $\texttt{Con}_{\texttt{spec}}$ specification at the cost of more complex concepts. We plan to explore this direction more in the future.
Defn. 2 describes the implementation that we use for the concept representation map $rep$ that uses a VLM $g$. This definition assumes that concepts are represented as directions in the VLM space; however note that this notion was enforced during training. In particular, the VLM’s training objective requires the pair of embeddings generated for an image-caption pair, by the image and text encoders of a VLM, to be aligned in the same direction as measured via cosine similarity.
@@
The direction corresponding to a concept $con$ is extracted using the text encoder $g_{enc}^{txt}$ of $g$. Following previous work [25], a vector $\overline{con}$ in the direction of concept is computed as the mean of text embeddings for a set of captions (denoted as $cap(con)$) that all refer to $con$ in different ways. We compute the mean since there may be many different valid captions constructed for the same concept and each caption leads to a slightly different embedding. Notice that there is no need for manual concept annotations to infer the concept directions.
@@
Given a VLM $g$ with image encoder $g_{enc}^{img}:X\rightarrow Z$ and text encoder $g_{enc}^{txt}:T\rightarrow Z$ where $Z:=\mathbb{R}^{d^{\prime}}$, the linear concept representation map, $rep$, via model $g$ is defined as,
@@
We can use Defn. 2 to provide an implementation for $rep$ which we statistically validate to check that it is semantically meaningful, according to human-understandable concepts (as described in Section 5.4). We can use $rep$ to instantiate the semantics of our language and reason about properties for the VLM (i.e., CLIP) itself. In principle, we can use the same implementation of $rep$ to reason about the original vision model. However, this would introduce a major scalability challenge—in the course of verifying a vision model $f$, we would not only need to reason about the whole model $f$ but also about the complex image encoder $g_{enc}^{img}$ of VLM $g$ since $g_{enc}^{img}$ is used to implement $rep$. Fortunately, it has been previously observed [25] that the representation spaces of vision-based DNNs and VLMs such as CLIP can be linked to each other via an affine map. We employ this map to define a new implementation of $rep$ which results in scalable verification.
@@
Let $D_{\text{train}}\subset X$ denote an image dataset. Given the encoder $f_{enc}:X\rightarrow Z_{f}$ of a vision model $f$ and the image encoder $g_{enc}^{img}:X\rightarrow Z_{g}$ of VLM $g$, we aim to learn a map $r_{map}:Z_{f}\rightarrow Z_{g}$ such that the representation spaces of the two models are aligned as much as possible, i.e., given an image $x\in X$, ideally $r_{map}(f_{enc}(x))=g_{enc}^{img}(x)$ holds.
Moayeri et al. [25] show that $r_{map}$ can be restricted to the class of affine transformations, i.e., $r_{map}(z):=Mz+d$, where $r_{map}$ is learnt by solving the following optimization problem,
@@
We use $r_{map}$ to define $rep$ as follows.
Definition 3 (Linear $rep$ via vision model, $r_{map}$, and VLM)
Given a vision model $f:X\rightarrow Y$ with encoder $f_{enc}:X\rightarrow Z_{f}$, a VLM $g$ with encoders $g_{enc}^{img}:X\rightarrow Z_{g}$ and $g_{enc}^{txt}:T\rightarrow Z_{g}$, and a linear map $r_{map}$ as described above, then the concept representation map, $rep$, is defined as,
@@
where $\overline{con}$ is a vector in $Z_{g}$ whose direction corresponds to concept $con$, i.e., $\overline{con}:=\frac{\Sigma_{t\in cap(con)}g_{enc}^{txt}(t)}{|cap(con)|}$.
We analyze some of the properties of this $rep$ as follows. First, let us formalize the notion of faithful representation space alignment (Defn. 4). Intuitively, the alignment is faithful if no information is lost when using $r_{map}$ to map between the two representation spaces.
Definition 4 (Faithful alignment of representation spaces)
$\forall x\in X.~{}r_{map}(f_{enc}(x))=g_{enc}^{img}(x)$
It is easy to see that if $r_{map}$ satisfies the condition of faithful alignment, then the implementations of $rep$ according to Defn. 2 and 3 coincide. Thus we could use Defn. 3 instead of Defn. 2 when implementing $rep$. However, in practice, it is possible that the representations spaces of a vision model and a VLM may not be faithfully aligned; there may be cases such that for the $r_{map}$ we learn, $r_{map}(f_{enc}(x))=g_{enc}^{img}(x)$ does not hold for some images. Thus, it is possible that the two implementations of $rep$ as we defined (one via VLM and one via vision model and $r_{map}$) may yield different results for some images. However, we argue that since we need to statistically validate $rep$ anyway, we can directly statistically validate $rep$ (from Defn. 3) and use it to instantiate the semantics of our language.
Once we have an implementation of $rep$ as per Defn. 3, we can define a constraint-based verification procedure for the vision model $f$. This implementation of $rep$ has the interesting property that it allows us to factor out the constraints corresponding to $f_{enc}$ and $g_{enc}^{img}$ resulting in efficient verification.
@@
We exploit the decomposition of a vision model $f$ into $f_{enc}$ and $f_{head}$ to define an efficient verification procedure for vision models with respect to a $\texttt{Con}_{\texttt{spec}}$ specification. In particular, Thm. 4.1 shows that the verification of a model $f$ with respect to a $\texttt{Con}_{\texttt{spec}}$ specification $e$ can be reduced to the verification of $f_{head}$, without involving $g_{enc}^{img}$.
@@
Given a vision model $f:X\rightarrow Y$ that can be decomposed into $f_{enc}:X\rightarrow Z$ and $f_{head}:Z\rightarrow Y$ where $Z:=\mathbb{R}^{d^{\prime}}$, a $\texttt{Con}_{\texttt{spec}}$ specification $e$, a linear concept representation map $rep$, as defined in Defn. 3, and an input scope $B\subseteq X$,
@@
where $\widehat{rep}$ operates on embeddings instead of inputs and is obtained from $rep$ by replacing $f_{enc}$ with the identity function and $f_{enc}(B):=\{f_{enc}(x)\mid x\in B\}$.
@@
See Appendix 0.A.
Discussion.
Our verification is sound only with respect to the interpretation given to the predicates by the concept representation map ($rep$). In this paper, we use a VLM to implement $rep$; VLMs are trained on massive amounts of data and, as a result, are a rich repository of concepts. With the trend towards training even bigger VLMs on larger amounts of data, the implementation of $rep$ is expected to improve. We use a statistical analysis to validate that $rep$ conforms with human understanding, i.e., it is semantically meaningful (described in Section 5.4). Further, we restrict our verification to concepts that are well represented in the VLM; if some concept is not well represented, we would not attempt verification for it. Note also that Thm. 4.1 holds for arbitrary decompositions of $f$ into $f_{head}$ and $f_{enc}$. Our assumption, supported by previous work [25], is that the head of the network is moderately sized, making it suitable for analysis with modern decision procedures. Although these are strong assumptions, they enable us to make progress on the challenging problem of verification of semantic properties for vision models.
@@ 5 Case Study
5.1 Dataset, Concepts, and Strength Predicates
@@
[24] is a dataset consisting of 26k images drawn from ImageNet [9], organized into 10 classes matching those of CIFAR10. The dataset contains manually annotated instance-wise labels for 18 attributes (which we use to define set $Concepts$) as well as the respective segmentation masks for these attributes on the images. The dataset is partitioned into a train set with 21k images and a test set with 5k images.
@@
The information in Figure 4 can be further used to elicit strength predicates relevant for each class $c$—for each pair of relevant ($con_{r}$) and irrelevant ($con_{ir}$) concepts, we can define a predicate $con_{r}\texttt{>}con_{ir}$ that we expect to hold on inputs of class $c$. For instance, for class truck, we formulate a total of 72 predicates, corresponding to all the possible combinations of relevant and irrelevant concepts; e.g., we elicit $wheels\texttt{>}wings$ and $metallic\texttt{>}tail$, as $wheels$ and $metallic$ are relevant concepts, whereas $wings$ and $tail$ are not. Overall, we formulate a total of 522 such predicates for the model. In practice, we expect human domain experts to directly provide the concepts and strength predicates relevant to the task. Since we only use the concept annotations to elicit the strength predicates, there is no need to manually annotate each image if these predicates are given.
@@ 5.3 Extraction of Concept Representations
In order to analyze the models with respect to the specifications that we formulate for RIVAL10, we need to build the concept representation map $rep$ using the VLM (as defined in Defn. 3). In particular, we need to extract the directions corresponding to the relevant concepts in CLIP’s representation space and learn the affine map from the representation space of ResNet18 to CLIP. The first step is to define the set of relevant concepts. For RIVAL10, these are the 18 attributes defined by the developers of the datatset. In general, the relevant concepts are elicited in collaboration with domain experts.
Next, to extract the directions corresponding to these 18 concepts, we use CLIP’s text encoder (referred to as $g_{enc}^{txt}$) in a manner similar to the approach described by Moayeri et al. [25]. In particular, as described in Defn. 2, for each concept $con$, we create a set of captions that refer to the concept (denoted as $cap(con)$). As an example, for the concept metallic, we use captions such as “a photo containing a metallic object”, “a photo of a metallic object”, etc. The complete set of captions used is given in Appendix 0.B. We then apply $g_{enc}^{txt}$ to each caption in the set $cap(con)$ and compute the mean of the resulting embeddings. The direction of the resulting mean vector corresponds to $\overline{con}$, i.e., the direction representing the concept $con$ in CLIP’s representation space.
[…]